OpenStudy (anonymous):

7+x-(3+32/8)=3 please show work.

OpenStudy (anonymous):

7+x-(3 + 32/8) = 3 32/8 is equal to 4 7 + x - (3+4)= 3 7 + x -7 = 3 so 7-7 is 0 x = 3